# https://github.com/czg1225/SlimSAM

### v3

@@ Abstract
Current approaches for compressing the Segment Anything Model (SAM) yield commendable results, yet necessitate extensive data to train a new network from scratch. Employing conventional pruning techniques can remarkably reduce data requirements but would suffer from a degradation in performance. To address this challenging trade-off, we introduce SlimSAM, a novel data-efficient SAM compression method that achieves superior performance with extremely less training data. The essence of SlimSAM is encapsulated in the alternate slimming framework which effectively enhances knowledge inheritance under severely limited training data availability and exceptional pruning ratio. Diverging from prior techniques, our framework progressively compresses the model by alternately pruning and distilling distinct, decoupled sub-structures. Disturbed Taylor pruning is also proposed to address the misalignment between the pruning objective and training target, thereby boosting the post-distillation after pruning. SlimSAM yields significant performance improvements while demanding over 10 times less training data than any other existing compression methods. Even when compared to the original SAM, SlimSAM achieves approaching performance while reducing parameter counts to merely 1.4% (9.1M), MACs to 0.8% (23G), and requiring only 0.1% (10k) of the SAM training data.
Keywords:
Segment Anything Model Compression Data-Efficient
@@ 1 Introduction
To mitigate these constraints, many efforts [21, 50, 48, 41, 51] have been made to effectively compress SAM. Without exception, these endeavors opt to replace the originally heavyweight image encoder with a lightweight and efficient architecture. This invariably entails training a new network from scratch. With regard to scratch training, an unavoidable challenging trade-off arises between training costs and model performance. As depicted in Figure 1, existing methods all inevitably compromise performance when training with very limited data.
@@
Figure 1: (a) In contrast to scratch training or distilling from randomly initialized networks, SlimSAM achieves superior performance with significantly reduced training data requirements. (b) Maintaining a similar compression ratio, SlimSAM also outperforms other competing methodologies.
@@
As depicted in Figure 1, comprehensive assessments across performance metrics, efficiency, and training data requirements reveal that SlimSAM markedly enhances compression performance, concurrently achieving superior lightweight and efficiency with markedly reduced training data requirements. Notably, our entire compression can be completed using only 10k un-labeled images on a single Titan RTX GPU within a span of merely 1 to 2 days.
@@ 3.1 Identifying SAM Redundancy
$I_{w_{i}}$ $\approx\left|\mathcal{L}_{w_{i}=0}(x_{i},y_{i})-\mathcal{L}_{w_{i}=0}(x_{i},y_{i})+\frac{\partial\mathcal{L}(x_{i},y_{i})}{\partial w_{i}}w_{i}\right|,$ (3)
$=\left|\frac{\partial\mathcal{L}(x_{i},y_{i})}{\partial w_{i}}w_{i}\right|.$
However, there exist two distinct limitations associated with the above Taylor importance estimation when pruning the image encoder of SAM. Firstly, the accuracy of Taylor importance relies heavily on the availability of sufficiently accurate hard labels $y_{i}$. Unfortunately, due to the intricate nature of jointly optimizing the image encoder and combined decoder [48], the post-distillation process necessitates performing on the image embedding $t_{i}$, resulting in the utilization of soft labels exclusively. Secondly, a concern arises regarding the consistency of loss functions when employing Taylor importance estimation for SAM pruning. The importance estimation strategy’s primary objective is to identify parameters $w_{i}$ that minimize the hard label discrepancy $|\Delta\mathcal{L}(x_{i},y_{i})|$. In contrast, the goal of the distillation-based recovery process is to minimize the soft label loss $|\Delta\mathcal{L}(x_{i},t_{i})|$. This misalignment in optimization objectives potentially impedes the efficacy of the distillation process. The experimental results in 5 also strongly prove our conclusion.
@@
Remark. The implementation of alternate slimming on decoupled sub-structures significantly reduces the disruption to the original model, particularly when the pruning ratio is quite high. This strategy also preserves consistent dimensionality, enabling effective intermediate feature distillation, especially beneficial in data-scarce conditions. Consequently, in comparison to the previous pruning and distillation methods, our alternate slimming achieves a 3.40% and 0.92% increase in MIoU when the pruning ratios achieve 77% and 50%.
@@ 4 Experiments
Method Paramsnormal-$\downarrow$ MACsnormal-$\downarrow$ TrainSet BatchSize GPUs Iters MIoUnormal-$\uparrow$
SAM-H [18]
641M
2736G
11M(100%)
256
256
90k
78.30%
SAM-L [18]
312M
1315G
11M(100%)
128
128
180k
77.67%
SAM-B [18]
93M
372G
11M(100%)
128
128
180k
73.37%
FastSAM-s [50]
11M
37G
220k(2%)
32
8
625K
30.72%
FastSAM-x [50]
68M
330G
220k(2%)
32
8
625K
35.41%
MobileSAM [48]
9.8M
40G
100k(1%)
8
1
100k
62.73%
EfficientSAM-t [41]
10M
28G
12.2M(110%)
128
64
450k
69.42%
EfficientSAM-s [41]
26M
94G
12.2M(110%)
128
64
450k
71.19%
EdgeSAM [51]
9.6M
23G
100k(1%)
64
8
50k
65.96%
@@
Ratio Method Labels Paramsnormal-$\downarrow$ MACsnormal-$\downarrow$ MIoUnormal-$\uparrow$
Ratio=0%
SAM-H [18]
641M
2736G
78.30%
SAM-L [18]
312M
1315G
77.67%
SAM-B [18]
93M
372G
73.37%
Ratio=50%
Scratch Distillation
26M 98G
1.63%
Random Pruning
71.03%
Magnitude Pruning [10]
69.96%
Hessian Pruning [22]
71.01%
Taylor Pruning [30]
71.15%
@@
Ratio=77%
Scratch Distillation
9.1M 23G
1.34%
Random Pruning
62.58%
Magnitude Pruning [10]
61.60%
Hessian Pruning [22]
63.56%
Taylor Pruning [30]
64.26%
@@ 4.2 Comparision and Analysis
Comparing with existing SAM compression methods. As depicted in Table 1, we conducted a comprehensive comparison encompassing performance, efficiency, and training costs with other SOTA methods. Our SlimSAM-50 and SlimSAM-77 models achieve a remarkable parameter reduction to only 4.0% (26M) and 1.4% (9.1M)of the original count, while also significantly lowering computational demands to just 3.5% (98G) and 0.8% (23G) MACs, all while maintaining performance levels comparable to the original SAM-H. In contrast to other compressed models, our approach yields substantial performance enhancements while simultaneously achieving greater lightweight and efficiency. SlimSAM consistently delivers more accurate and detailed segmentation results across various prompts, preserving SAM’s robust segmentation capabilities to the greatest extent. This qualitative superiority over other models is visually evident in Figure 4 and 5. Our approach demonstrates outstanding levels of accuracy and correctness. Most notably, SlimSAM achieves these remarkable outcomes with exceptionally low training data requirements, utilizing merely 0.1% (10k) images of the SA-1B dataset. This represents a significant reduction in data dependency, requiring 10 times less data than both EdgeSAM and MobileSAM, and 1,100 times less data than EfficientSAM.
Figure 5: Comparison of segment everything results. In comparison to other models, our segmentation results exhibit greater proximity to the original SAM-H in terms of both accuracy and comprehensiveness.
@@ 5 Ablation Study and Analysis
Table 3: Comparison between proposed disturbed Taylor pruning and original Taylor pruning.
@@
Method MIoUnormal-$\uparrow$
Taylor Pruning
62.04%
Disturbed Taylor Pruning 62.31%
SlimSAM-77 + Taylor
63.63%
SlimSAM-77 + Disturbed Taylor 64.48%
Step Distillation Objective MIoUnormal-$\uparrow$
Step 1
Final Image Embeddings
65.10%
@@
Step 2
Final Image Embeddings
63.91%
@@
Table 6: Comparision of training results using varied amounts of training data but maintaining the same iterations.
Method Normalization MIoUnormal-$\uparrow$
Local Pruning
64.38%
Global Pruning
Mean
63.64%
Max
64.35%
Sum
63.55%
@@
Standardization
64.14%
Pruning Ratio Data Iters MIoUnormal-$\uparrow$
@@
5k
200k
64.47%
2k
200k
61.72%
Table 6: Comparision of training results using varied amounts of training data but maintaining the same iterations.
@@
Global Pruning vs Local Pruning. Finally, we conducted experiments to evaluate the performance of local pruning and global pruning in bottleneck pruning. Given that the bottleneck dimensions in each block are entirely decoupled, we systematically applied channel-wise group pruning at various ratios for each block while preserving the predefined overall pruning ratio in this step. To obtain a consistent global ranking, we normalized the group importance scores $I_{\mathcal{G}}$ of each layer in five ways: (i) Sum: $I_{\mathcal{G}i}=\frac{I_{\mathcal{G}i}}{\sum_{i=1}^{K}I_{\mathcal{G}i}}$, (ii) Mean: $I_{\mathcal{G}i}=\frac{I_{\mathcal{G}i}}{\sum_{i=1}^{K}I_{\mathcal{G}i}/K}$, (iii) Max: $I_{\mathcal{G}i}=\frac{I_{\mathcal{G}i}}{Max_{i=1}^{K}(I_{\mathcal{G}i})}$, (iv) Standarization: $I_{\mathcal{G}i}=\frac{I_{\mathcal{G}i}-Max_{i=1}^{K}(I_{\mathcal{G}i})}{Max_{i=1}^{K}(I_{\mathcal{G}i})-Min_{i=1}^{K}(I_{\mathcal{G}i})+1e-8}$, (v) Gaussian: $I_{\mathcal{G}i}=\frac{I_{\mathcal{G}i}-\sum_{i=1}^{K}I_{\mathcal{G}i}/K}{\sigma_{i=1}^{K}(I_{\mathcal{G}i})+1e-8}$. As indicated in Table 6, local pruning ensures consistent performance, whereas global pruning raises the model’s upper-performance limit. Global pruning’s efficacy is highly dependent on the chosen importance normalization method. For our model, we opted for global pruning with Gaussian normalization, which yielded the best training results. Following global pruning, Figure 7 illustrates the dimensions of bottleneck features (QKV embeddings and MLP hidden embeddings) within each ViT in the image encoder. When applying mean, sum, or Gaussian normalization, the ViTs in the middle exhibit more group redundancy compared to those at the beginning and end. However, the pruned dimensions do not display distinct patterns when utilizing max or standardization normalization.

### v4

# 0.1% Data Makes Segment Anything Slim

@@ Abstract
Current approaches for compressing the Segment Anything Model (SAM) yield commendable results, yet necessitate extensive data to train a new network from scratch. Employing conventional pruning techniques can remarkably reduce data requirements but would suffer from a degradation in performance. To address this challenging trade-off, we introduce SlimSAM, a novel data-efficient SAM compression method that achieves superior performance with extremely less training data. The essence of SlimSAM is encapsulated in the alternate slimming framework which effectively enhances knowledge inheritance under severely limited training data availability and exceptional pruning ratio. Diverging from prior techniques, our framework progressively compresses the model by alternately pruning and distilling distinct, decoupled sub-structures. Disturbed Taylor pruning is also proposed to address the misalignment between the pruning objective and training target, thereby boosting the post-distillation after pruning. SlimSAM yields significant performance improvements while demanding over 10 times less training data than any other existing compression methods. Even when compared to the original SAM, SlimSAM achieves approaching performance while reducing parameter counts to merely 1.4% (9.1M), MACs to 0.8% (23G), and requiring only 0.1% (10k) of the SAM training data. Code is available at https://github.com/czg1225/SlimSAM
@@ 1 Introduction
To mitigate these constraints, many efforts [21, 50, 48, 41, 51] have been made to effectively compress SAM. Without exception, these endeavors opt to replace the originally heavyweight image encoder with a lightweight and efficient architecture. This invariably entails training a new network from scratch. With regard to scratch training, an unavoidable challenging trade-off arises between training costs and model performance. Existing methods all inevitably compromise performance when training with very limited data.
@@
Comprehensive assessments across performance metrics, efficiency, and training data requirements reveal that SlimSAM markedly enhances compression performance, concurrently achieving superior lightweight and efficiency with markedly reduced training data requirements. Notably, our entire compression can be completed using only 10k un-labeled images on a single Titan RTX GPU.
@@ 3.1 Identifying SAM Redundancy
$I_{w_{i}}$ $\approx\left|\mathcal{L}_{w_{i}=0}(x_{i},y_{i})-\mathcal{L}_{w_{i}=0}(x_{i},y_{i})+\frac{\partial\mathcal{L}(x_{i},y_{i})}{\partial w_{i}}w_{i}\right|=\left|\frac{\partial\mathcal{L}(x_{i},y_{i})}{\partial w_{i}}w_{i}\right|.$ (3)
However, there exist two distinct limitations associated with the above Taylor importance estimation when pruning the image encoder of SAM. Firstly, the accuracy of Taylor importance relies heavily on the availability of sufficiently accurate hard labels $y_{i}$. Unfortunately, due to the intricate nature of jointly optimizing the image encoder and combined decoder [48], the post-distillation process necessitates performing on the image embedding $t_{i}$, resulting in the utilization of soft labels exclusively. Secondly, a concern arises regarding the consistency of loss functions when employing Taylor importance estimation for SAM pruning. The importance estimation strategy’s primary objective is to identify parameters $w_{i}$ that minimize the hard label discrepancy $|\Delta\mathcal{L}(x_{i},y_{i})|$. In contrast, the goal of the distillation-based recovery process is to minimize the soft label loss $|\Delta\mathcal{L}(x_{i},t_{i})|$. This misalignment in optimization objectives potentially impedes the efficacy of the distillation process. The experimental results in Section 5 also strongly prove our conclusion.
@@
Remark. The implementation of alternate slimming on decoupled sub-structures significantly reduces the disruption to the original model, particularly when the pruning ratio is quite high. This strategy also preserves consistent dimensionality, enabling effective intermediate feature distillation, which is especially beneficial in data-scarce conditions. Consequently, in comparison to the previous pruning and distillation methods, our alternate slimming achieves a 3.40% and 0.92% increase in MIoU when the pruning ratios achieve 77% and 50%.
@@ 4.1 Experimental Settings
[…]
SAM-H [18] 641M 2736G 11M(100%) 256 256 90k 78.30%
SAM-L [18] 312M 1315G 11M(100%) 128 128 180k 77.67%
SAM-B [18] 93M 372G 11M(100%) 128 128 180k 73.37%
FastSAM-s [50] 11M 37G 220k(2%) 32 8 625K 30.72%
FastSAM-x [50] 68M 330G 220k(2%) 32 8 625K 35.41%
MobileSAM [48] 9.8M 40G 100k(1%) 8 1 100k 62.73%
EfficientSAM-t [41] 10M 28G 12.2M(110%) 128 64 450k 69.42%
EfficientSAM-s [41] 26M 94G 12.2M(110%) 128 64 450k 71.19%
EdgeSAM [51] 9.6M 23G 100k(1%) 64 8 50k 65.96%
@@
[…]
Ratio=0% SAM-H [18] 641M 2736G 78.30%
SAM-L [18] 312M 1315G 77.67%
SAM-B [18] 93M 372G 73.37%
Ratio=50% Scratch Distillation 26M 98G 1.63%
Random Pruning 71.03%
Magnitude Pruning [10] 69.96%
Hessian Pruning [22] 71.01%
Taylor Pruning [30] 71.15%
@@
Ratio=77% Scratch Distillation 9.1M 23G 1.34%
Random Pruning 62.58%
Magnitude Pruning [10] 61.60%
Hessian Pruning [22] 63.56%
Taylor Pruning [30] 64.26%
@@ 4.2 Comparision and Analysis
Comparing with existing SAM compression methods. As depicted in Table 1, we conducted a comprehensive comparison encompassing performance, efficiency, and training costs with other SOTA methods. Our SlimSAM-50 and SlimSAM-77 models achieve a remarkable parameter reduction to only 4.0% (26M) and 1.4% (9.1M)of the original count, while also significantly lowering computational demands to just 3.5% (98G) and 0.8% (23G) MACs, all while maintaining performance levels comparable to the original SAM-H. In contrast to other compressed models, our approach yields substantial performance enhancements while simultaneously achieving greater lightweight and efficiency. SlimSAM consistently delivers more accurate and detailed segmentation results across various prompts, preserving SAM’s robust segmentation capabilities to the greatest extent. This qualitative superiority over other models is visually evident in Appendix Figure 8 and 7. Our approach demonstrates outstanding levels of accuracy and correctness. Most notably, SlimSAM achieves these remarkable outcomes with exceptionally low training data requirements, utilizing merely 0.1% (10k) images of the SA-1B dataset. This represents a significant reduction in data dependency, requiring 10 times less data than both EdgeSAM and MobileSAM, and 1,100 times less data than EfficientSAM.
@@ 5 Ablation Study and Analysis
Table 3: Comparison between disturbed Taylor pruning and original Taylor pruning.
[…]
Taylor Pruning 62.04%
Disturbed Taylor Pruning 62.31%
SlimSAM-77 + Taylor 63.63%
SlimSAM-77 + Disturbed Taylor 64.48%
@@
[…]
Step 1 Final Image Embeddings 65.10%
@@
Step 2 Final Image Embeddings 63.91%
@@
[…]
Local Pruning 64.38%
Global Pruning Mean 63.64%
Max 64.35%
Sum 63.55%
@@
Standardization 64.14%
Table 6: Comparision of training results using varied amounts of training data.
[…]
@@
5k 200k 64.47%
2k 200k 61.72%
@@
Global Pruning vs Local Pruning. Finally, we conducted experiments to evaluate the performance of local pruning and global pruning in bottleneck pruning. Given that the bottleneck dimensions in each block are entirely decoupled, we systematically applied channel-wise group pruning at various ratios for each block while preserving the predefined overall pruning ratio in this step. To obtain a consistent global ranking, we normalized the group importance scores $I_{\mathcal{G}}$ of each layer in five ways: (i) Sum: $I_{\mathcal{G}i}=\frac{I_{\mathcal{G}i}}{\sum_{i=1}^{K}I_{\mathcal{G}i}}$, (ii) Mean: $I_{\mathcal{G}i}=\frac{I_{\mathcal{G}i}}{\sum_{i=1}^{K}I_{\mathcal{G}i}/K}$, (iii) Max: $I_{\mathcal{G}i}=\frac{I_{\mathcal{G}i}}{Max_{i=1}^{K}(I_{\mathcal{G}i})}$, (iv) Standarization: $I_{\mathcal{G}i}=\frac{I_{\mathcal{G}i}-Max_{i=1}^{K}(I_{\mathcal{G}i})}{Max_{i=1}^{K}(I_{\mathcal{G}i})-Min_{i=1}^{K}(I_{\mathcal{G}i})+1e-8}$, (v) Gaussian: $I_{\mathcal{G}i}=\frac{I_{\mathcal{G}i}-\sum_{i=1}^{K}I_{\mathcal{G}i}/K}{\sigma_{i=1}^{K}(I_{\mathcal{G}i})+1e-8}$. As indicated in Table 6, local pruning ensures consistent performance, whereas global pruning raises the model’s upper-performance limit. Global pruning’s efficacy is highly dependent on the chosen importance normalization method. For our model, we opted for global pruning with Gaussian normalization, which yielded the best training results. Following global pruning, Figure 4 illustrates the dimensions of bottleneck features (QKV embeddings and MLP hidden embeddings) within each ViT in the image encoder. When applying mean, sum, or Gaussian normalization, the ViTs in the middle exhibit more group redundancy compared to those at the beginning and end. However, the pruned dimensions do not display distinct patterns when utilizing max or standardization normalization. The impact of global pruning becomes more pronounced with an increased number of training iterations. Specifically, when training extends to 80 epochs, the MIoU for global pruning exceeds that of local pruning by approximately 2%.
@@ 6 Conclusion
Appendix A Appendix
❈ In this document, we provide supplementary materials that extend beyond the scope of the main manuscript, constrained by space limitations. These additional materials include in-depth information about the ablation study of SlimSAM-50, further experiments assessing model efficiency, additional evaluations of training costs, analysis of dynamic loss, limitation discussion, and supplementary qualitative results.
A.1 Ablation Study on SlimSAM-50
We performed an extensive series of ablation studies on the SA-1B [18] dataset utilizing the SlimSAM-50 model, characterized by its significant 50% pruning ratio. To guarantee a fair and consistent comparison across these ablation studies, each model under evaluation was uniformly trained over a span of 20 epochs, employing a training dataset comprising 10,000 images.
Disturbed Taylor Pruning. Initially, we executed an ablation study to evaluate the effects of our innovative disturbed Taylor pruning technique on distillation processes. This approach strategically aligns pruning criteria with the optimization goals of the ensuing distillation, thereby facilitating enhanced performance recovery. As illustrated in Table 8, our disturbed Taylor pruning method consistently outperforms, achieving markedly better results at equivalent training expenditures. In comparison to the conventional one-step pruning strategy and our alternative slimming approach, our methodology registers MIoU enhancements of 0.31% and 0.39% over the standard Taylor pruning method [30], respectively.
Intermediate Aligning. We further investigated the impact of integrating alignment with intermediate layers in the distillation process. As Table 8 illustrates, leveraging knowledge from these intermediate layers substantially enhances the training outcomes. Specifically, when distillation in step 1 incorporates learning from both bottleneck features and final image embeddings, there is a notable 1.19% improvement in MIoU compared to a methodology reliant solely on image embeddings. Similarly, in step 2 of the distillation process, a strategy that utilizes both embedding features and final image embeddings demonstrates a 0.29% MIoU improvement over approaches exclusively based on image embeddings.
Table 7: Comparison between disturbed Taylor pruning and original Taylor pruning.
[…]
Taylor Pruning 70.02%
Disturbed Taylor Pruning 70.33%
SlimSAM-50 + Taylor 70.42%
SlimSAM-50 + Disturbed Taylor 70.81%
[…]
Step 1 Final Image Embeddings 70.86%
Step 1 + Bottleneck Features 72.07%
Step 2 Final Image Embeddings 70.52%
Step 2 + Embedding Features 70.81%
[…]
Local Pruning 70.81%
Global Pruning Mean 70.76%
Max 70.77%
Sum 70.80%
Gaussian 70.83%
Standardization 70.78%
Figure 5: The intermediate dimensions of QKV Attention (top row) and MLP (bottom row) within each ViT after pruning. We present the outcomes of local pruning and global pruning under five distinct normalization methods.
Figure 6: Training results on SA-1B with common one-step strategy and our alternate slimming strategy.
Alternate Slimming. Moreover, we conducted a series of experiments to examine the efficacy of our novel alternate slimming strategy. Diverging from the traditional one-step pruning approach, our method divides structural pruning into two distinct and progressive phases. In the initial phase, pruning is exclusively focused on the dimensions pertaining to embedding features. The subsequent stage then targets dimensions associated with bottleneck features. After completing both embedding and bottleneck pruning, we employ knowledge distillation with intermediate layer alignment on the pruned model to facilitate performance restoration. The outcomes derived from our proposed disturbed Taylor importance estimation are displayed in Figure 6. This figure demonstrates that our alternative slimming strategy significantly boosts MIoU, achieving an increase of 0.5%. When juxtaposed with the ablation study results of SlimSAM-77, it becomes evident that our strategy exhibits a more pronounced improvement, particularly when applied to models with higher pruning ratios.
Global Pruning vs Local Pruning. In our final set of experiments, we assessed the effectiveness of both local and global pruning approaches in the context of bottleneck pruning. Considering the complete decoupling of bottleneck dimensions in each block, we meticulously implemented channel-wise group pruning at varying ratios across different blocks. This was done while maintaining the predetermined overall pruning ratio for this phase of the study. To obtain a consistent global ranking, we normalized the group importance scores $I_{\mathcal{G}}$ of each layer in five ways: (i) Sum: $I_{\mathcal{G}i}=\frac{I_{\mathcal{G}i}}{\sum_{i=1}^{K}I_{\mathcal{G}i}}$, (ii) Mean: $I_{\mathcal{G}i}=\frac{I_{\mathcal{G}i}}{\sum_{i=1}^{K}I_{\mathcal{G}i}/K}$, (iii) Max: $I_{\mathcal{G}i}=\frac{I_{\mathcal{G}i}}{Max_{i=1}^{K}(I_{\mathcal{G}i})}$, (iv) Standarization: $I_{\mathcal{G}i}=\frac{I_{\mathcal{G}i}-Max_{i=1}^{K}(I_{\mathcal{G}i})}{Max_{i=1}^{K}(I_{\mathcal{G}i})-Min_{i=1}^{K}(I_{\mathcal{G}i})+1e-8}$, (v) Gaussian: $I_{\mathcal{G}i}=\frac{I_{\mathcal{G}i}-\sum_{i=1}^{K}I_{\mathcal{G}i}/K}{\sigma_{i=1}^{K}(I_{\mathcal{G}i})+1e-8}$. Table 9 reveals that while local pruning maintains consistent performance, global pruning enhances the upper limit of the model’s capabilities. In a departure from the findings observed with SlimSAM-77, various normalization methods do not markedly influence post-pruning performance. This suggests that the necessity of selecting an optimal normalization technique increases with the pruning ratio. For our model, we chose global pruning combined with Gaussian normalization, which led to the most favorable training outcomes. Figure 5 delineates the distribution of bottleneck feature dimensions (including QKV and MLP hidden embeddings) across each Vision Transformer (ViT) [6] in the image encoder. When mean, sum or Gaussian normalization is applied, the dimensions within the ViTs tend to distribute more evenly. However, employing max or standardization normalization often results in significant variances in the intermediate dimensions of each ViT.
A.2 More Analysis on Efficiency
In the principal manuscript, the high efficiency of our SlimSAM model is objectively substantiated through the disclosure of parameter counts and Multiply-Accumulate Operations (MACs). This section extends the evaluation by reporting on actual acceleration in inference, further affirming the model’s efficiency. As delineated in Table 10, SlimSAM-50 outperforms the original SAM-H model by achieving a 6.9-fold increase in inference speed, while SlimSAM-77 attains an 8.3-fold acceleration. Our compression methodology markedly diminishes the actual inference time, concurrently effecting substantial reductions in both model size and MACs. The inference acceleration metrics were tested using an NVIDIA TITAN RTX GPU.
Table 10: Inference acceleration was empirically tested on an NVIDIA TITAN RTX GPU, revealing that higher pruning rates significantly improve inference speed.
[…]
Ratio=0% SAM-H [18] Faster×1.0
SAM-L [18] Faster×1.7
SAM-B [18] Faster×4.3
Ratio=50% SlimSAM-50(Ours) Faster×6.9
Ratio=77% SlimSAM-77(Ours) Faster×8.6
Table 11: The training results were compared using the varied amounts of training data but maintaining the same training iterations.
[…]
Ratio=50% 10k 100k 72.33%
5k 100k 71.89%
2k 100k 69.79%
Ratio=77% 10k 200k 67.40%
5k 200k 64.47%
2k 200k 61.72%
A.3 More Analysis on Training Costs
In our foundational manuscripts, we demonstrate that SlimSAM exhibits exceptional compression performance with minimal training cost. A pertinent inquiry emerges: can SlimSAM maintain its competitive performance with reduced training costs? To address this, we have undertaken supplementary experiments focusing on the interplay between training cost and performance.
In Table 11, we present the results of additional experiments conducted with varying quantities of training data, while keeping the number of training iterations constant. We observe that with a pruning ratio of 50%, a reduction in the volume of training data only marginally impacts the model’s performance. Notably, even when trained with a limited dataset of just 2,000 images, our SlimSAM-50 model remarkably attains an MIoU of nearly 70%. However, as the pruning ratio is elevated to 77%, a decrease in training data more significantly affects performance. This leads to the inference that although our methodology, which integrates pruning and distillation techniques, mitigates the need for extensive training datasets, the availability of more training data can still enhance model performance, particularly at higher pruning rates. It can be anticipated that with an increase in the volume of training data, our model may potentially achieve lossless compression of SAM.
Table 12 showcases the outcomes of experiments conducted by modifying the parameters for training iterations, while maintaining a constant training dataset size. The results clearly illustrate a direct relationship between the quantity of training iterations and the effectiveness of the model compression. It is evident that more extensive training significantly improves the performance of our compressed models. Remarkably, SlimSAM maintains its superiority over other methods even when the training iterations are halved, demonstrating its robustness and efficiency in achieving high-performance compression.
Table 12: Training outcomes were evaluated using the same amount of training data across different numbers of training iterations.
[…]
Ratio=50% 10k 50k 70.83%
10k 100k 72.33%
Ratio=77% 10k 100k 64.43%
10k 200k 67.40%
Table 13: Ablation study on dynamic loss weights for distillation
Step Model Constant ($\alpha=0.5$) Dynamic ($\alpha=\frac{N-n-1}{N}$)
1 SlimSAM-50 MIoU:72.07% MIoU:71.60%
SlimSAM-77 MIoU:66.32% MIoU:65.79%
2 SlimSAM-50 MIoU:70.42% MIoU:70.83%
SlimSAM-77 MIoU:63.63% MIoU:64.48%
A.4 More Analysis on Dynamic Loss
Further experiments were undertaken to assess the efficacy of employing dynamic loss within our intermediate feature alignment procedure. The outcomes of these ablation studies are detailed in Table 13. It was observed that a constant weight mechanism is more apt for scenarios involving a robust teacher model, whereas the implementation of a dynamic weight strategy enhances performance in instances where the teacher model exhibits lesser strength.
A.5 Limitations
In this analysis, we critically examine the constraints of our methodology.
First, our approach demonstrates robust compression performance with minimal training data. Nonetheless, an expanded training dataset could further enhance the model’s capabilities. Our current pre-trained SlimSAMs, limited by hardware constraints, are trained on a dataset of only 10,000 images from the SA-1B dataset. Utilizing a more comprehensive training dataset could potentially enable our method to achieve lossless compression.
Second, the essence of our method lies in employing structural pruning and knowledge distillation to preserve the knowledge of original pre-trained SAMs. This strategy inherently sets the performance ceiling of our model at the level of the original SAM. We found it challenging to surpass the performance of the original SAM, which acted as both the target for pruning and the target for optimization. A key area for future research will be exploring how to surpass the performance of the original SAM with limited parameter counts and reduced training costs.
A.6 More Qualitative Results
We present more visual comparisons with other existing compressed models and the original SAM-H. Figure 8, 7 and 9 provide a detailed visual comparison using the segment-everything prompt, while Figures 10 and 11 showcase additional qualitative results obtained with box prompts and point prompts, respectively. Relative to established compression models such as MobileSAM [48], FastSAM [50], EdgeSAM [51] and EfficientSAM [41], our model distinctly outperforms in achieving more precise segmentation, particularly noticeable at the object edges. Notably, even when benchmarked against SAM-H, our model demonstrates commensurate segmentation capabilities.
A.7 Societal impacts
In this paper, we introduce SlimSAM, a novel data-efficient SAM compression method that delivers superior performance using minimal training data. SlimSAM achieves an outstanding compression ratio while preserving robust segmentation capabilities. This advancement enables the deployment of SAM on resource-constrained edge devices, underscoring its significant practical applications.
Figure 7: Comparison of segmentation results using segment everything prompts
Figure 9: Comparison of segmentation results using segment everything prompts
Figure 10: Comparison of segmentation results using box prompts
Figure 11: Comparison of segmentation results using point prompts